09/21/05 -This project is being moved to the archive because the whole project is being rebid due to bidding anomalies.

09/15/05 – After much discussion amongst the vendor, the contractor, Tony and Amy in purchasing, it was decided that TE did not use proper bidding procedure on this job and it has to be reibd. TE will provide the new bid documents and prebid and contract time at no cost to the County. TE still has to write an engineering change order for the scope change for the new electrical design (TA to talk to Tony about this new SOSS). The bid and the construction admin will seem to be au gratis from this point forward.

09/07/05 – ThermalTech has received many calls from the vendors on this job mostly in regards to the chiller selection on the project. Trane believes the chiller must be individually test to meet ARI standards and York seems to think that the specifications say that only so many out of 100 need to be test to meet ARI. TTC assured Tony that ThermalTech will review all aspects of the schedule, specifications and bids and provide a written recommendation as to which chiller meets the intent of the specification while maintaining the requirements of the ORC. TTC will ask Tony about getting the purchasing department or prosecutor department to review and provide a decision if needed after the contract is approved. As with any public bid, the bid amount cannot change so the vendors will have to provide the equipment listed to the County at the bid amount or other actions will have to be taken.

09/01/05 – Contracts came back from Prosecutor. TTC to pick up, modify and take back to faculties for approval. Peck and Trane and calling multiple times trying to force us to use the Trane chiller because it is the only to meet the ARI specification. TTC to review ARI 550 and will discuss with everyone after the contracts are signed..

08/24/05 – TTC, TA, AKM and JM talked about how to write the Admin Chiller Contracts and what alternates to take based on our latest discussions with Cinergy and what they are supplying to the building. TTC has the contracts written to accept the EC base bid and Alternate #1 (new service) and for the MC we are taking base bid and Alternate #3 (warranty). TTC stated that the contracts will be dropped off today.

08/08/05 – TE and the County had a meeting today to discuss how to write contracts for this project and what alternates to accept. For the mechanical contractor we will accept base bid and Alternate #3 for 5 year parts and labor warranty. In discussing the scope of the electrical alternates, it was discovered that to accept the new feeder to the 10th floor does not truly provide the full extent of the County desire to have backup power from both grids to building at all times. It is a convoluted situation but basically the building now has 8000 amps to the building from 2 separate grid feeds. If either grid goes down, the building can still maintain power from the other grid. The max usage right now is 3700 on 95 degree OA temperatures. If we only feed the new chiller with the new service, the main building will have no additional power savings. In the end it was decided that the best thing to do was to try to have the contractor provide a new 1600 amp service into the basement from both grids and provide a transfer switch in the basement as well. This will slightly modify the original scope of the Alternate but will handle TE’s misunderstanding of what the County requested in the design process of this job and in no way would have changed the outcome of the bid process. TE will research the changes and advise the County on possible cost and acceptance by Cinergy.

08/01/05 – The bids opened last Thursday. TE is working on which recommendations to accept for this project. The numbers varied for using bus duct in lieu of standard conduit and feeder from a $400 deduct to a $127,000 adder. Because of this TE has to do some research to see what is correct. TE believes that we will be able to accept Alternate #1, Alternate #3 and Alternate #4 and still be under budget for this job. We want to write the contract on either Tuesday or Wednesday of this week.

07/20/05 – Tim Centers from Habegger called to ask about freeze protection on the chiller since we aren’t using glycol at the Admin building. TZ is on vacation but we are using heat tape and DDC alarms to protect the chiller and piping. Ryan Halverson from Trane called to ask about removing the word “Scroll” from the schedule so everyone is bidding “Screw” compressor chillers and TTC will do that in the Addendum.

07/18/05 – TE to work on Addendum 01. Fix the bond issues for purchasing. Move rigging time to full weekend. Move conduit pull to weekend. Add01 should go out on Tuesday.

07/12/05 – The bids are out and the prebid is set for 07/15/05 at 1:30 p.m.

06/15/05 – JM reported that discreet chiller control will be fine for this job.

06/09/05 – TE emailed Tony and Joe about using discreet points to run the chiller plant instead of adding a Trace interface. TE needs to wait for approval from Tony

05/20/05 – TTC dropped off the chiller plans for review. The specs are not complete yet for review and TTC has not decided the best way to bid the chiller (LCC, Alternates or Base and Equals). GD has returned his comments to TTC.

04/27/05 – TTC, TA, TZ, Joe Merkt and AKM met to discuss the project so that TE can design in a way that doesn’t go too far over budget. The cost estimate turned in last year only had $33,000 for electrical upgrade. The plan is to make or use the mechanical scope adjustable so that some additional monies can be diverted to upgrade the electrical. TE has to do more fieldwork to check on the piping arrangement so that TE can fully comprehend the additional return pipe that goes from the 10th floor to the basement. Basically the 3-way valves have a return system that goes from floor to floor UP the building and from the 10th floor back down to the basement.

04/22/05 – This week THP has delivered final structural dwgs to TE for project. TE finalized a control strategy within TE for the new chiller and discussed hydraulic issues of existing piping and pumps. GD gave TZ and TTC direction on sizing the new 10th floor pump. TTC decided that JCI controls can be modified during the upcoming project and that existing control work on 10th floor chiller would have no adverse effects on new system. Chiller S/S and Pump S/S and statuses can be re-used and extended as needed during 10th floor project. Trane sent project info to TE on the existing chillers: Evaporator, 45-55 F, 480 gpm, 2-pass, 26.5 ft wpd, .0005 fouling factor, Condenser: 85-95 F, 581 gpm, 2-pass, 21.7 ft wpd, .0005 fouling factor, CAB-CH-1 TRANE CENTRAVAC CVHE-020B L85B26377, CAB-CH-2 TRANE CENTRAVAC CVHE-020B L85B26376.

04/05/05 – TE coordinated with THP on steel locations and scope for new air cooled chiller. THP will meet onsite with Joe Merkt to field check existing steel conditions this week. THP requested $2000 for fee’s for this project. TE approved. This cost wasn’t in the original scope of fee’s but TE will try to incorporate into project but will save for engineering change order if needed.

03/21/05 – TE kicked off project with a site visit for TA, TTC and TZ. SRB visited the site for equipment and piping takeoffs during the same morning. TE began talking to vendors about equipment weights and chiller selections. TE discussed the cost estimate and various upgrade options at the building for this project.

02/09/05 – Received Signed SOSS

01/12/05 – Project Kickoff meeting with County. TE to begin design.

Project Status

03/24/05 – TTC sent the revised change order to ICS for him to sign and prepared it for mailing, stopped them and reprinted them again because Gary called Nancy to schedule the City and Fire Marshal inspection to find that cost is $475 additional to the city fire inspection so now I am advocating that the change order be re-written back up to $1400 to cover the additional charge and Gary will just absorb the rest of the overage from his profits. Larry Keifer has required a FULL BUILDING HORN & STROBE test. Joe called TTC to express his disbelief that a full building test would be required and indicated how difficult that would be to do in one hour. TTC sent an email to Larry Keifer requesting clarification and hopefully explanation on this upcoming test. Tentatively the test is scheduled for Tuesday March 29 at 6:00 a.m.
03/23/05 – ICS finished relocating the fire alarm strobe for our new fire alarm permit. Joe tested the strobe and it works. TTC talked to ICS about painting the wall where the old strobe was moved and negotiated a $450 reduction in the original change order down to $950 and in return the County can touch up the column at their convenience. TTC and JM agreed that this seemed reasonable. TE will rewrite the Change Order and get ICS to sign it for final billing. ICS has to coordinate a final inspection between Larry Keifer and Nancy at the Cincinnati Fire Dept. He hopes it will happen this week, but TE believes this will take at least a week to get scheduled. ICS is to coordinate with Joe Merkt when FA test is actually scheduled.

03/15/05 – ICS submitted a change order price of $1400 to move the strobes for the approved fire alarm permit. Diana approved. TE will now scheduled (we have already called to schedule the work) and assist in the installation of strobes as needed.

03/08/05 – City Permit Office reported today that Fire has approved and they are in the process of setting up the permit. They will call me at TE when ready.

03/07/05 – TTC called the City Permit desk today. Plan Examiner approved on 02/22/05. The Fire Department has the permit now for approval. TTC called 352-4314 and convinced them that the Fire Sprinkler work was completed under Permit # 2004P06759. The Fire Inspector updated the database. Now the city should call us in a couple of days with the final permit price and tell me it is ready to pick up. TTC will call ICS to get them to pay for the work and give us a CO price to do the work for this job.

03/03/05 – TTC contacted the city permit department about status – no word yet.

02/22/05 – TTC contacted the city about permit status – no word received back at this time. We will continue to push for final inspections. TTC authorized ICS to bill for $3500 leaving a $200 balance on the project and whatever this change order turns out to be since all of the original work is complete and punched.

02/14/05 – TTC requested final paperwork from Carrigan again. TTC requested allowance usages from B&J on project for final bookkeeping.

02/11/05 – TTC contacted Preferred about lowering the sprinkler head in the public work area. ICS called to report that the final IBI inspection was today and that everything passed. Final paperwork is en route. Preferred called back and will get it taken care of quickly through Joe Merkt.

02/11/05 – TE submitted the Fire Alarm Permit in person today. The FA permit number is 2005P01862. It will have to be paid for when it is picked up. The review should take under 12 days. TTC asked for Bill Spurling to do the review. TE will keep everyone apprised of progress.

02/10/05 – TE has finished recreated 4 sets of new fire alarm permits for the microfilm area. MS talked to the building department today and we can try for a mini permit but the drawings have to be on 8.5×11’s only. These plans are on 11×17 and stamped and signed so we are going to submit them as a standard FA set and not walk them through as that doubles the cost. ICS will have to move one strobe still to meet the new plans we are submitting.

01/31/05 – TTC talked to Bill Spurling at City plans exam. He did not know where Larry wanted the additional strobe – perhaps the darkroom. He asked me to talk to Larry, so I placed a call to Mr. Keifer and when he calls me back, I will get to the bottom of what we need to do and get it taken care of immediately.

01/27/05 – TE mailed Gary @ Integrated blank final document forms: G706, G706A, G707, Affidavit of Contractor Prevailing Wages and, Affidavit of Original or Sub-Contractor.

01/25/05 – City is now requiring a possible new strobe in the public area which would mean we have to take out a fire alarm permit. Stiles still believes that you can get a TCO for the space and we can pull that permit and get that work done and still make the city happy. TE has placed two calls into William Spurling, the city Plans Examiner, to discuss the issue with him. TE will keep the County apprised.

01/21/05 – Project passed above ceiling inspection with IBI. EC is 95% complete. Larry Keifer may still require a Fire Alarm permit to be pulled for this job although TE doesn’t believe that is necessary. TE coordinate with Preferred to add a sprinkler head in the vestibule as it was shown on the drawings but not to add the one above the microfiche charts as it would tend to get damaged from the anticipated use plus there is not 18″ clearance above the racks.

01/18/05 – Plumbing is finished including vent. HVAC has installed toggle switch for exhaust fan. Auxiliary exhaust fan is burned out so TE instruct B&J to buy a new motor and replace it. B&J has the motor in stock now and will install it on Wed Jan 19th.

01/12/05 – Dennis Grimm of Carrigan & Grimm contacted, requesting 3 additional documents (Affidavit of Original or Subcontractor, Affidavit of Contractor Compliance-Prevailing Wages and Consent of Surety to Final Payment) to complete submission of all final documents. Dennis not sure how to obtain the forms therefore the forms were mailed to Carrigan & Grimm on this date.

12/27/04 – TE contacted ICS about catching up with GC. They said they would be on site this evening at 4:00pm to catch up. ICS will also begin installed lead anchor bolts with 1/4″ #20 eyelets for Communications to hang their J-Hooks for data cable support from. TE advised them not to exceed $640 dollars for this work. Gary Kercher agreed to this pricing scheme. DES can contact Mitch that they will be able to move their cables to the new supports soon as this will need to be done before final inspection.

12/10/04 – TE advised the GBBN can create a wooden soffit around conduit and panel box in vestibule and paint the wood and the cover of the panel and still meet code requirements for the panel. County arranged to have carpenters do the work through a work order. Matthew shall enter a work order into the system.

12/07/04 – Construction Mtg addressed the scheduled as we are about 1 week behind due to holidays and with Christmas approaching, the project is likely to stay a week behind. County asked TE to look at cleaning up existing panel in vestibule for better aesthetics. PC is complete. HVAC will be in on Thursday night to relocated controller for hood fan.

12/01/04 – GBBN & TE resolved the RFI requested by ICS electric in regards to the existing electric panel in Room #201 near the countertop. Microfilm has agreed to redesign the countertop to accommodate the electric panel. GBBN will follow through with 3R and TE will notify ICS that no additional work will be required on their part.

11/24/04 – Rough-in inspection complete. Inspector may request clean up of existing cabling above ceiling. TE to investigate.

11/03/04 – Asbestos Memo was issued to contractors for tile removal by pinnacle communication. PDF. Preferred (Jeff) is confirmed for evening work tonight (11/03/04) and we have coordinated with Joe Merkt for this evening work. TE has contacted all contractors about not working on 11-04-04.

11/01/04 – Carrigan called today to discuss existing piping layout, mixing valves, and hot water heater placement. It appears that some ball valves were added to self regulate temperature at the film equipment. After discussing the situation with Dennis, we believe it would be better to pipe the building hot water into the water heater and have it “boost” that water up to inlet hot water temperature at the mixing valves thereby giving the process room immediate hot water when needed. In it current configuration it backfeeds the building hot water system which would give us unpredictable results. TE will confirm this at tomorrows meeting but in the meantime, we have asked Carrigan to give us a price to replace both temperature mixing valves with new Simmons Temptrol mixing valves that have integral temperature sensors for visual tracking outlet temperatures and piping the water heater in series with the buildings hot water supply to the process area. We will discuss all this at tomorrow’s meeting.

10/27/04 – Fire Permit has been approved: 2004P06759. Carrigan & Grimm schedule of values received and approved.

10/26/04 – Code questions were received by permit office. ThermalTech responded with these answers. PDF

10/25/04 – Carrigan asked to rework the vent arrangement on the sink and equipment in the new area to prevent a flat vent condition on the floor drain. TE has approved this routing of the new vent. Minor change and not cost to the project.

10/19/04 – Const Mtg 2 held. Plumbers, HVAC, Electricians, Fire Protection on board and ready to begin evening work. Scope coordinated through GBBN during project meeting. Owner agreed to allow ceilings to be removed during remodel work and replaced at very end as long as contractors take out the minimum required for each phase.

10/07/04 – Const Mtg held. Carrigan & Grimm was not present. All work appears ready to start on October 18. Phasing and project schedule mostly approved. Preferred needs security checks approved and DES will send them to JJ. We may see additional monies requested from EC for additional permit cost. We may or may not approve these additional funds depending on how the project goes. We can use allowance money if needed and nothing else comes up. TE will post minutes after they are received from GBBN.